OpenStudy (anonymous):

Let f(x)=2x and g(x) = x^2 +4. Find (g • f)(-3)

OpenStudy (akashdeepdeb):

g(x) = \(x^2 + 4\) f(x) = 2\(x\) (g o f)(x) = g ( f(x) ) Thus, (g o f)(-3) = g ( f(-3) ) Now calculate f(-3). Let f(-3) = a Your answer then will be g(a). Got it?

OpenStudy (anonymous):

So would f(-3) be 2(-3)?

OpenStudy (math&ing001):

Yes @Dreaaaaaa

OpenStudy (math&ing001):

Can you carry on ?

OpenStudy (anonymous):

So then the answer for f(x) would be -6 and I have to plug that into g(x)?

OpenStudy (math&ing001):

Exactly !

OpenStudy (anonymous):

Okay thank you! I understand now !

OpenStudy (math&ing001):

You're welcome, have a great time at Openstudy !

OpenStudy (precal):

is this a composition? I believe it is times

OpenStudy (anonymous):

What do you mean?

OpenStudy (precal):

is your dot a "dot product" which means to multiply or it is a hollow dot which means composition. It does matter because the other responses are for composition but I think it it solid and therefore it means multiplication

OpenStudy (anonymous):

No it's not a solid it's an open circle. I just didn't have it on my keyboard :/

OpenStudy (precal):

ok then an open circle is composition and the above instructions are correct.